Resources for Dealing with Encryption Problems

March 28, 2001

1 Conversion between text and numbers

Note that some schemes involve converting sequences of, perhaps ten, characters simultaneously to numbers. One method, of course, is manual table lookup. A table for conversion to ASCII is available with a simple encryption example^{1}. Conversion between text and numbers can be done in many ways with a computer program written in almost any programming language.

Conversion of text to ASCII decimal Example of a program in the language “C” for converting text to ASCII: anum.c.

Conversion of ASCII decimal to text Example of a program in the language “C” for converting ASCII to text: numa.c.

2 Encryption of sequences of numbers

Maple code Code for computing a high power of a sequence of numbers for a given modulus: hipo.

Sample Maple session A sample Maple session for encoding and then decoding the ASCII decimal sequence for the word “Tomorrow”: msession.

1URI: ./adoc.html